OpenStudy (anonymous):

If <1 is congruent to <2, then

OpenStudy (anonymous):

http://tinypic.com/r/n2lso2/6 1. 3 is congruent to 4 2. <3 and <4 are complementary 3. ray BD is perpendicular line AC 4. <5 and is congruent to <8 5. segment AD is congruent to segment DC

OpenStudy (anonymous):

pic is not available

jimthompson5910 (jim_thompson5910):

If <1 is congruent to <2, and <1 and <2 are a linear pair, then x+x = 180 2x = 180 x = 180/2 x = 90 So angles 1 and 2 are both equal to 90 degrees

jimthompson5910 (jim_thompson5910):

Probably more useful is just the fact that angle 1 is 90 degrees you can use this to get your answer